The width of a rectangle is 7 meters less than the length. The perimeter is 66 meters. Find the length and the width.
Thanks so much.

Answer by TimothyLamb(4379)

w = L - 7
2L + 2w = 66
---
2L + 2(L - 7) = 66
2L + 2L - 14 = 66
4L = 66 + 14
L = (66 + 14)/4
L = 20
---
w = L - 7
w = 20 - 7
w = 13
---
L = 20 m
w = 13 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php